OpenStudy (anonymous):

Find the product of z1 and z2 where z1 = 8(cos 40° + i sin 40°), and z2 = 4(cos 135° + i sin 135°). OPTIONS: 32(cos 40° + i sin 40°) 12(cos 175° + i sin 175°) 32(cos 5400° + i sin 5400°) 32(cos 175° + i sin 175°) PLEASE HELP PLEASE

jimthompson5910 (jim_thompson5910):

Hint: p*cis(x) times q*cis(y) = (p*q)*cis(x+y) where cis(x) = cos(x) + i*sin(x)

OpenStudy (anonymous):

I dont get it :(

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

z1*z2 [8(cos 40° + i sin 40°)] * [4(cos 135° + i sin 135°)] [8*cis(40)] * [4*cis(135)] does that help?

OpenStudy (anonymous):

so its A right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

z1*z2 [8(cos 40° + i sin 40°)] * [4(cos 135° + i sin 135°)] [8*cis(40)] * [4*cis(135)] (8*4)*cis(40+135) ... use the formula I gave you See the answer now?

OpenStudy (anonymous):

Its D then @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

One more ?

OpenStudy (anonymous):

Find the cube roots of 27(cos 279° + i sin 279°). Could you help with this @jim_thompson5910

jimthompson5910 (jim_thompson5910):

use the formula found on this page http://orion.math.iastate.edu/trig/sp/xm08/applets/complexcuberoot.html

OpenStudy (anonymous):

AHH @jim_thompson5910 CONFUSION

jimthompson5910 (jim_thompson5910):

all you're doing is plugging in k = 0, k = 1, k = 2 to find the 3 cube roots

jimthompson5910 (jim_thompson5910):

r = 27 theta = 279

jimthompson5910 (jim_thompson5910):

hopefully this isn't brand new to you

OpenStudy (anonymous):

let me try

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

^zk= ^3sqrt 27(cos (279 +2kpi/3) + isin (279+2kpi/3)) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

now to find the 3 cube roots, you substitute k = 0, k = 1, and k = 3 you do each value of k one at a time

OpenStudy (anonymous):

That I dont know how to do

jimthompson5910 (jim_thompson5910):

so for example, when k = 0, the first cube root is z0 = 3*(cos(279+2*0*pi/3) + i*sin(279+2*0*pi/3)) z0 = 3*(cos(279/3) + i*sin(279/3)) z0 = 3*(cos(93) + i*sin(93)) So that is your first cube root

jimthompson5910 (jim_thompson5910):

plug in k = 1 to find the next cube root

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

wouldn't it be 3(-sin(279+pi/6) + icos(279+pi/6)

OpenStudy (anonymous):

? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

close z1 = 3*(cos(279+2*1*pi/3) + i*sin(279+2*1*pi/3)) z1 = 3*(cos( 279+2pi/3) + i*sin(279+2pi/3))

jimthompson5910 (jim_thompson5910):

and I'd just leave it like that

OpenStudy (anonymous):

And the last let me try

OpenStudy (anonymous):

eek

OpenStudy (anonymous):

is it just 3(cos (279) + isin (279)) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

unfortunately it's not that clean cut

OpenStudy (anonymous):

:(

jimthompson5910 (jim_thompson5910):

you plugged in k = 2 right?

OpenStudy (anonymous):

3 right ?

OpenStudy (anonymous):

Im confused about this one help:( sorry

jimthompson5910 (jim_thompson5910):

no k = 2

OpenStudy (anonymous):

I thought I only am supposed to plug in k = 0, k = 1, and k = 3

jimthompson5910 (jim_thompson5910):

no, k = 0, k = 1, k = 2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

like the page describes

jimthompson5910 (jim_thompson5910):

z2 = 3*(cos(279+2*2*pi/3) + i*sin(279+2*2*pi/3)) z2 = 3*(cos(279+4pi/3) + i*sin(279+4pi/3)) and i'd leave it like that

OpenStudy (anonymous):

THATS IT I had a different answer THANK YOU SO MUCH is there anything else ?

jimthompson5910 (jim_thompson5910):

nope, we found all 3 roots using the correct formula

jimthompson5910 (jim_thompson5910):

so we're done

OpenStudy (anonymous):

There is your medal as promised :)

OpenStudy (anonymous):

can i ask what happened to the 27 in the problem? it was in the formula then it seemed like it disappeared .. @jim_thompson5910